OpenStudy (anonymous):

Factor the polynomial completely. That is, once you are finished factoring, none of the factors you obtain should be factorable: 100x^2 − 180xy + 81y^2

OpenStudy (anonymous):

ok, help me work it out.

OpenStudy (anonymous):

Do you know how to find the GCF? (Greatest Common Factor)

OpenStudy (anonymous):

Check this out.. might help you. http://www.purplemath.com/modules/simpfact.htm

OpenStudy (anonymous):

sometimes lol uhh, lets see...this one is hard, kinda.

OpenStudy (anonymous):

10! i think

OpenStudy (anonymous):

100x^(2)-180xy+81y^(2) For a polynomial of the form ax^(2)+bx+c, find two factors of a*c (8100) that add up to b (-180).In this problem -(9)/(10)*-(9)/(10)=(81)/(100) (which is (c)/(a)) and -(9)/(10)-(9)/(10)=-(180)/(100) (which is ((b)/(a)) , so insert -(9)/(10) as the right hand term of one factor and -(9)/(10) as the right-hand term of the other factor. (x-(9)/(10))(x-(9)/(10)) Remove the fraction by multiplying the first term of the factor by the denominator of the second term. (10x-9y)(10x-9y) Combine the two common factors of (10x-9y) by adding the exponents. (10x-9y)^(1+1) Add 1 to 1 to get 2. (10x-9y)^(2)

OpenStudy (anonymous):

I was suppose to do this by myself lol

OpenStudy (anonymous):

The final answer would be (10x−9y)2

OpenStudy (anonymous):

Oh pssh.. I was going to put dont look till you finished :p..

OpenStudy (anonymous):

lmaoo too late! :P lol NEXT!

OpenStudy (anonymous):

xDD